Shift Summary for evening shift, 30/july/99

S. Elizabeth Turner (eturner@kudzu)
Fri, 30 Jul 1999 23:58:36 -0500

Shift Summary for evening shift, 30/july/99

On Shift: BobT,SET

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: ComEd shutdown, problems bringing
accelerator back up after shutdown
========================================================================
Runs ended this shift:
none
=======================================================================
Tapes brought to FCC this shift:
(including tapes for online split)
none
========================================================================
Accesses during shift:

BobT and SET make two accesses to reinstall HA S/S power supply.
BobT and DougJ make an access to replace PC terminal monitor.
========================================================================
Overview of shift:
No beam.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
16:00 SET and BobT installed HA S/S power supply with
no load.
17:25 DougJ and BobT replaced bad PC monitor and took
zpos readings.
18:20 L2 Busy -- reloaded Busy logic and L2 state
machine to fix.
19:30 DN put 3 new bottles of argon ethane on line
21:25 Ramped magnet up, then down to zero to prepare
for muon run. (Ramped up bc field was in a
'bad state' and the fix is to ramp up and then
back down).
21:50 SET and BobT put load of 4 S/S on HA S/S power
supply. So far so good. Note that there are
no signals going thru the S/S -- we are just
going to see if the power supply can handle the
load for a while.
========================================================================
Follow up Notes on Previous Shift

SDAQ button on Etot-temp is still red. The read out is 0.69 V instead
of 0.80 (corresponding to 80 degrees). Jason H verified this downstairs
on the digital thermometers. It doesn't seem to be any worries for now.
The red button can be ignored.

========================================================================
Items to Follow up on Next Shift

SDAQ S9-UCLA-PS4 button is no longer red. If it turns red, it is not an
emergency. SET would like to be notified but not awakened.

Take a muon run when beam returns. Then run at negative polarity.